QUESTION 6 Calculate AG (in kJ) at 35.4°C for a reaction that has keq = 124
Answer #1

T= 35.4 oC

= (35.4+273.1) K

= 308.5 K

use:

ΔGo = -R*T*ln Kc

= - 8.314*308.5* ln(1.24*10^2)

= -12359 J

= -12.4 KJ

Answer: -12.4 KJ